# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### 12 Mar 2026 v3

@@ 17 Conclusion: The Solution of Planar QCD
In this paper, we have presented the exact analytic solution of the planar MM loop equations, thereby solving large-$N_{c}$ QCD in the continuum limit. Building on Part I, where the confining dressing factor $\exp\left(-\kappa S[C]\right)$ was shown to arise from the rigid Hodge-dual additive minimal surface, we have constructed the complete dynamics without summing over fluctuating worldsheet metrics.
Our first main outcome is structural: by working natively in momentum loop space, we bypass the coordinate-space contact singularities and cusp pathologies that obscure the local continuum limit of the MM loop equations. We rigorously established that a purely one-dimensional analytic Taylor-Magnus solution of the exact vector momentum-loop equation cannot exist. The expansion is mathematically consistent through $\mathcal{O}(P^{\prime 6})$ but violently breaks down at $\mathcal{W}^{(8)}$ due to a Fredholm rank deficiency. This topological obstruction proves that a 1D closed-loop scalar framework lacks the internal degrees of freedom to absorb the macroscopic spatial stress of Planar QCD. This anomaly natively forces the introduction of 4D continuous boundary-twistor variables to parametrize the rigid minimal surface, yielding the confining twistor-string representation.
Our second main outcome is dynamical: we demonstrated that the internal Majorana fermions (the “elves”) on the worldsheet provide the precise algebraic mechanism required for planar factorization at intersections. Crucially, the bulk Liouville term in our effective action originates microscopically from the zero-point fluctuations of these hidden fermions. The conformal anomaly of these heavy elves yields exactly the Lüscher term ($\pi/12$) in the effective string potential, cleanly generating the universally correct intercepts 1/96-1/96 and $47/96$ for the physical $\pi$ and $\rho$ Regge trajectories without any ad hoc assumptions of macroscopic string vibrations.
One may view this construction as a geometric realization of the Master Field [54]—the existence of a purely classical description for the large-$N_{c}$ limit of QCD. While this Master Field does not manifest as a single classical gauge field configuration in coordinate spacetime, it finds an exact geometric representation in loop space. The classical geometry uncovered here identifies the underlying master object as the rigid Hodge-dual minimal surface. Consequently, the phase-space path integral over quark loops, once transformed into two-dimensional boundary twistor variables, exhibits exact geometric localization governed by Catastrophe Theory. The true Master Field of Planar QCD is thus a critical classical trajectory in twistor space—a geometric structure conceptually parallel to the spectral curves of Seiberg-Witten theory.
By establishing the Master Field as this rigid twistor geometry, our approach naturally avoids the “graviton problem” typical of AdS/CFT gravity holography. Because our bulk geometry is not dynamical, there are no propagating bulk graviton modes that would lack a physical counterpart in confining QCD. All propagating degrees of freedom are strictly confined to the boundary, while bulk quantities (such as the Liouville field) are rigidly determined by the analytic continuation of boundary data. Therefore, our framework is best understood as gauge holography, where the bulk serves as a rigid encoding of the boundary dynamics rather than gravity holography or a topological string theory.
Ultimately, Planar QCD ceases to be a theory of stochastic quantum fluctuations; it becomes a theory of pure classical complex geometry. The extraction of the exact mass spectrum reduces from a fluctuating path integral to a deterministic Generalized Eigenvalue Problem. By summing over the purely imaginary monodromy phases of the twistor poles protected by the pinch-singularity boundary barrier, we obtain an exact Bohr-Sommerfeld quantization condition.
For the fundamental one-pole trajectory, this yields the unified parametric Regge trajectory (1.1) with trigonometric functions of the twistor branch point degree $a$, and $q=0$ for the unshifted unnatural parity ($\pi$, $K$) trajectories, while q=1q=-1 for the shifted natural parity ($\rho$) trajectory. Although these functions are mathematically non-linear, they are numerically remarkably close to linear Regge trajectories. This analytic solution passes through the experimental data on Figure 15. The only outlier ($3\sigma$) is $K_{3}(2320)$, which is marked as “unconfirmed” in the PDG tables. The exact mass spectrum in our solution of Planar QCD is thus deterministically governed by the degenerate saddle points of the complexified twistor-string.
17.1 Future Directions: Glueballs, Baryons, and Scattering
The representation of Planar QCD as an Analytic Twistor String opens several immediate avenues for future research:
  • 1.
    Glueballs: The entire construction, from the loop equations and the fermion determinant on the minimal surface to the localization of the twistor path integral, naturally generalizes to higher $1/N_{c}$ corrections. Resolving the mathematical specifics of this expansion is the natural next step to analytically extract the glueball spectrum.
  • 2.
    Baryons: In the large-$N_{c}$ limit, baryons appear as heavy states composed of $N_{c}$ quarks. Geometrically, this corresponds to a minimal surface with a specific topology (e.g., a “Y-junction” or a surface with multiple boundaries). Extending the twistor ansatz to surfaces with higher connectivity will allow for the exact geometric quantization of baryon masses.
  • 3.
    Scattering Amplitudes: Meson-meson scattering involves a minimal surface bounded by four segments. The crossing symmetry and Regge behavior of these amplitudes should emerge naturally from the analytic continuation of the twistor data, potentially connecting this confining geometric framework directly to modern Amplituhedron methods.
@@ Acknowledgements
I wish to express my deepest gratitude to Nima Arkani-Hamed for his hospitality and for the invitation to present the complete derivation of this solution in a series of lectures at the IAS Particle Physics (Pizza) Seminar. The intellectual rigor and stimulating environment he fosters at the Institute were instrumental in the final crystallization of these results.
I am especially indebted to Edward Witten for his profound insights and for the invaluable discussions that clarified the correspondence between this twistor-geometric solution and the phenomenological QCD string observed in lattice simulations. His generous engagement with the topological aspects of this work has been a source of significant inspiration.
@@
During the preparation of this work, the author utilized the generative AI Gemini 3 Flash to assist in the technical refinement of the manuscript. Specifically, the tool was employed to:
  • 1.
    Identify and correct typographical errors and grammatical inconsistencies.
  • 2.
    Standardize mathematical notation and ensure the rigorous formatting of LaTeX equations across the text and appendices.
  • 3.
    Improve the clarity and stylistic flow of the scientific narrative while preserving the original physical derivation.
  • 4.
    Cross-verify the consistency of the unified mass-squared formula with experimental data fits and topological sector identifications.
The author has thoroughly reviewed and edited all AI-generated refinements and takes full responsibility for the physical content, mathematical accuracy, and original conclusions presented in this work.
@@ Appendix C Explicit invariant tensor solutions for the MLE
Table 1: Linear algebra diagnostics for the exact Vector Momentum Loop Equation. Here $\Delta\text{rank}:=\text{rank}([A_{n}|b_{n}])-\text{rank}(A_{n})$. A strict positivity, $\Delta\text{rank}>0$, rigorously proves the system is mathematically inconsistent.
@@ C.1 Solving for the lowest terms $W^{(n<8)}$
The tensor $\mathcal{W}^{(6)}_{\mu_{1}\dots\mu_{6}}$ is completely determined by the geometric stress cascading from the left-hand side loop derivative acting on the lower-order terms, successfully absorbing the kinematic constraints without contradiction. Its explicit expansion in terms of the 15 independent 6-point Kronecker pairings is uniquely fixed by the lower-order parameters $c_{2,1}\equiv 1$, $c_{4,1}$, and $c_{4,2}$.
At $\mathcal{W}^{(8)}$, this linear absorption fails. The expansion truncates due to the irreducible non-linear $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$ stress, which exceeds the tensor rank of the pure cyclic Shuffle Ideal basis, rigorously proving the non-analyticity of the functional $\mathcal{W}[P]$.
@@ C.2 Counting of unknowns/equations and Fredholm inconsistency at $\mathcal{W}^{(8)}$
To rigorously demonstrate the breakdown of the Taylor-Magnus expansion at the 8th order ($\mathcal{W}^{(8)}$), we cast the Momentum Loop Equation at each order $n$ into a finite linear system. After imposing the closed-loop shuffle-ideal reduction (loop-closing prescription), cyclic symmetry, and parity constraints (all odd orders vanish), the exact functional matching reduces to:
@@
The total number of unknown variables solved for at order $n$ is therefore $N_{unk}(n)=N_{\mathcal{W}}(n)+N_{free}(<n)+N_{gauge}(n)$. We analyzed the exact linear algebra of these systems using a specialized symbolic script. The exact matrix ranks and subsystem counts are summarized in Table 1.
@@
At $n=8$, the matrix $A_{8}$ possesses 297 left-null vectors. Projecting the inhomogeneous right-hand side $b_{8}$ (the geometric stress generated by the discrete non-linear cross-terms of $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$) onto these null directions yields explicit, non-zero rational obstructions. For example, selecting the first canonical left-null basis vector $u_{1}$ yields the exact rational projection:
@@
This establishes a decisive mathematical contradiction. Two critical robustness statements follow directly from this matrix representation:
@@
The full basis construction, shuffle-ideal reduction, and exact linear-algebra verification—including the generation of the matrices $A_{n}$ and the explicit null-vector projections—are provided in the executable Mathematica notebook [34] (see Data Availability).
Appendix D Chiral Twistors and the Minkowski Helicoid Metric
In this appendix, we explicitly derive the macroscopic worldsheet metric of the rotating string directly from the microscopic twistor fields. We demonstrate how the chiral twistor parameterization breaks the dimensional reduction to a 3-dimensional hyperplane, and how the physical parity-even projection of the double Wick rotation analytically continues the Euclidean twistor metric into the physical Minkowski Nambu-Goto metric. We then pedantically evaluate the resulting bulk action integral.
D.1 Chiral Twistors and Dimensional Breakout
The canonical Helicoid is generated by twistors endowed with fractional powers of the worldsheet coordinate $z$. This fractional scaling is a spin-structure gauge choice that explicitly enforces the Neveu-Schwarz anti-periodic boundary conditions for the worldsheet fermions as they complete a full rotation around the string center. To construct the chiral (Hodge-dual) surface, we set the right twistor to a constant basis spinor. We define the left and right twistors as:
$\lambda(z)=e^{\mathrm{i}\pi/4}z^{-1/2}\begin{pmatrix}1/z\\ 1\end{pmatrix}$,
$\mu(z)=\mathrm{i}e^{-\mathrm{i}\pi/4}z^{1/2}\begin{pmatrix}1\\ 0\end{pmatrix}$ (D.1)
Because $\lambda\propto z^{-1/2}$ and $\mu\propto z^{1/2}$, their product has no overall fractional branching. The resulting bosonic physical observables are strictly rational, making this configuration geometrically equivalent to a rational single-pole state.
The defining geometric characteristic of this configuration is its behavior under the symmetric Pauli matrices. The spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$. Evaluating the symmetric bilinear for the first spatial component:
$f^{\prime}_{1}\propto\lambda^{T}\sigma_{1}\mu=\lambda_{1}\mu_{2}+\lambda_{2}\mu_{1}$
$=\left(e^{\mathrm{i}\pi/4}z^{-3/2}\right)(0)+\left(e^{\mathrm{i}\pi/4}z^{-1/2}\right)\left(\mathrm{i}e^{-\mathrm{i}\pi/4}z^{1/2}\right)=\mathrm{i}$ (D.2)
Unlike the generic configuration where this component vanished identically everywhere on the worldsheet, here the tangent vector component $f^{\prime}_{1}$ evaluates to a non-zero constant. Consequently, the target space coordinate $X_{1}$ is no longer strictly constant. This permanently breaks the exact algebraic constraint that previously crushed the generic minimal surface into a 3-dimensional hyperplane, meaning this chiral surface genuinely spans a 4-dimensional embedding.
D.2 The Euclidean Metric and Parity Projection
To extract the physical string metric, we compute the induced conformal factor $\exp\left(2\rho\right)=|\lambda|^{2}|\mu|^{2}$. Restoring the dimensionful scale parameter $R$, the squared norms are:
$|\lambda|^{2}=\frac{R}{|z|}\left(\frac{1}{|z|^{2}}+1\right)=R\frac{1+|z|^{2}}{|z|^{3}},$
$|\mu|^{2}=R|z|(1+0)=R|z|$ (D.3)
When multiplied, the fractional $|z|$ weights cancel perfectly:
$\exp\left(2\rho\right)=R^{2}\frac{1+|z|^{2}}{|z|^{2}}=R^{2}\left(1+\frac{1}{|z|^{2}}\right)$ (D.4)
To map this conformal geometry to the natural coordinates of the string worldsheet, we transition from the complex plane to the Euclidean cylinder via $z=\exp\left(\sigma_{E}+\mathrm{i}\tau_{E}\right)$. The flat line element transforms as $|dz|^{2}=|z|^{2}(d\sigma_{E}^{2}+d\tau_{E}^{2})=e^{2\sigma_{E}}(d\sigma_{E}^{2}+d\tau_{E}^{2})$. Substituting $|z|=\exp\left(\sigma_{E}\right)$, the conformal factor becomes:
$\exp\left(2\rho\right)=R^{2}\left(1+e^{-2\sigma_{E}}\right)$ (D.5)
Thus, the induced Euclidean worldsheet metric is exactly:
$ds^{2}_{E}=\exp\left(2\rho\right)e^{2\sigma_{E}}(d\sigma_{E}^{2}+d\tau_{E}^{2})$
$=R^{2}\left(1+e^{2\sigma_{E}}\right)\left(d\sigma_{E}^{2}+d\tau_{E}^{2}\right)$ (D.6)
To evaluate the physical area, we analytically continue this metric into Minkowski spacetime via a double Wick rotation. The angular parameter $\tau_{E}$ serves as the target space time. We perform a Wick rotation on the spatial worldsheet coordinate into the physical light-cone regime:
$\sigma_{E}=\mathrm{i}\sigma_{M}\implies d\sigma_{E}^{2}=-d\sigma_{M}^{2}$ (D.7)
Under this transformation, the exponential strictly rotates into a complex phase: $e^{2\mathrm{i}\sigma_{M}}=\cos(2\sigma_{M})+\mathrm{i}\sin(2\sigma_{M})$.
Because the macroscopic QCD string is strictly parity-conserving, the physical area is evaluated as the parity-even arithmetic mean of the left-chiral and right-chiral target space configurations. The conjugate right-chiral surface identically provides the inverse complex phase $e^{-2\mathrm{i}\sigma_{M}}$. Taking their arithmetic mean rigorously annihilates the imaginary sine terms, strictly isolating the real, physical part of the area:
$\text{Re}\left[\exp\left(2\rho\right)e^{2\mathrm{i}\sigma_{M}}\right]$ $=R^{2}\big(1+\cos(2\sigma_{M})\big)$
$=2R^{2}\cos^{2}(\sigma_{M})$ (D.8)
D.3 Action Evaluation and the Halved Lüscher Shift
Finally, we pedantically evaluate the total effective action over the fundamental domain of this purely real Minkowski metric. The determinant of the conformal Minkowski metric $ds^{2}_{M}=2R^{2}\cos^{2}(\sigma_{M})(d\tau_{E}^{2}-d\sigma_{M}^{2})$ yields the exact proper area measure:
$\sqrt{-g}=2R^{2}\cos^{2}(\sigma_{M})$ (D.9)
The center of the rotating string lies at $r=0$ ($\sigma_{M}=0$), and the physical speed-of-light boundary is at $r=1$ ($\sigma_{M}=\pi/2$). For one full spatial period of the meson, the time coordinate wraps exactly one cycle: $\tau_{E}\in[0,2\pi]$. The exact classical Minkowski area integral is therefore:
Area $=\int_{0}^{2\pi}d\tau_{E}\int_{0}^{\pi/2}2R^{2}\cos^{2}(\sigma_{M})d\sigma_{M}$
$=(2\pi)(2R^{2})\left(\frac{\pi}{4}\right)=\pi^{2}R^{2}$ (D.10)
Simultaneously, the Liouville action dictates that the worldsheet carries the kinetic anomaly $\frac{1}{12\pi}|\nabla\rho|^{2}$. Because the parity-symmetric conformal factor shares the exact same spatial singularity at the twistor pole as the generic string, it mathematically merges into the classical area term with the identical anomaly coefficient. As established in the main text, it acts strictly as a rigid, local shift to the effective string tension:
$\sigma_{\text{eff}}=\sigma+\frac{1}{12\pi R^{2}}$ (D.11)
Multiplying the purely real classical Minkowski area by this effective tension, we obtain the complete one-cycle bulk action:
$S^{(1)}_{bulk}$ $=-\sigma_{\text{eff}}\times\text{Area}$
$=-\left(\sigma+\frac{1}{12\pi R^{2}}\right)\left(\pi^{2}R^{2}\right)$
$=-\pi^{2}\sigma R^{2}-\frac{\pi}{12}$ (D.12)
This rigorous derivation explicitly demonstrates that for the parity-projected chiral string, the classical string area is exactly halved, which identically matches the universally mandated topological Lüscher phase $\pi/12$, though from a completely different physical and mathematical origins, which we discuss in the text.
Appendix E The Pinch Singularity and the Topological Barrier

### 23 Mar 2026 v4

@@ 17 Conclusion: A Solution of Planar QCD
In this paper, we have presented an analytic solution of the planar MM loop equations in the continuum limit. Building on Part I, where the confining dressing factor $\exp\left(-\kappa S[C]\right)$ arises from the Hodge-dual additive minimal surface, we have constructed the dynamics without summing over fluctuating worldsheet metrics.
Our first result is structural. In momentum loop space, the coordinate-space contact singularities and cusp pathologies are absent. The Taylor–Magnus expansion of the vector momentum-loop equation is consistent up to $O(P^{\prime 6})$ but develops a rank deficiency at $W(8)$. This algebraic obstruction shows that a one-dimensional scalar loop functional does not contain sufficient degrees of freedom to represent a solution of the vector loop equations. This necessitates the introduction of four-dimensional boundary twistor variables, leading to the twistor-string representation of the Hodge-dual minimal surface.
Our second result is dynamical. The internal Majorana fermions (the “elves”) on the Hodge-dual minimal surface provide the mechanism for planar factorization at intersections, in agreement with the right-hand side of the MM equations. The bulk Liouville term in the effective action arises from the cancelation of their zero-point fluctuations. This cancelation is needed to remove anomalies in the MM equations. The associated conformal anomaly yields the Lüscher phase ($\pi/12$) in the effective string action, contributing to the low-$M^{2}$ curvature of the Regge trajectories.
Within the planar limit ($N_{c}\to\infty$), the solution corresponds to a common confining string tension $\sigma$ for all trajectories. There are nonlinear effects, which are not present in the Nambu-Goto string theory, but these nonlinear effects are numerically small. The observed $\sim 10\%$ variation between the macroscopic slopes of the different flavor trajectories is attributed to $O(1/N_{c}^{2})$ corrections, including finite resonance widths, string interactions, and sea-quark effects, consistent with the expected accuracy of the planar approximation. The resulting parametric Regge trajectories describe thirteen meson families spanning the light, strange, charm, and bottom flavors ($\pi,\rho,K,K^{*},D,D^{*},D_{s},D_{s}^{*},B,B^{*},B_{s},B_{s}^{*},B_{c}$), reproducing the observed mass-spin relations across all 36 states (see Figure 15 and Table 1).
We must add a remark regarding the physical boundaries of the current geometric framework. The spontaneous breaking of chiral symmetry and the vanishing of the pion mass in the strict chiral limit are not explained by our theory. As demonstrated, the geometric pressure of the Liouville anomaly bounds the dimensionless string parameter to $x\geq 1/\sqrt{3\pi}$, which generates a strictly positive minimum geometric mass threshold ($2m\approx 270$ MeV) for the $J=0$ ground state. The physical $140$ MeV pion falls below this threshold because its anomalously light mass is governed by the non-perturbative chiral condensate, a phenomenon distinct from the geometry of the confining flux tube. Fully capturing the pseudo-Goldstone nature of the pion would require a structural modification of the theory, making the Master Field explicitly accommodate and respect spontaneous chiral symmetry breaking. We view the synthesis of this geometric confinement mechanism with chiral symmetry breaking as a crucial direction for future theoretical development.
The mathematical origin of our nonlinear Regge trajectories is radically different from that of conventional String Theory. The generation of S-matrix mass poles is associated with the degeneracies of the twistor-string effective action. After analytic continuation to Minkowski space, these poles correspond to energies at which the stability matrix develops a null space, producing non-compact directions in the associated Lefschetz thimbles.
These non-compact directions arise from the multi-valued monodromy of the holomorphic twistor fields. At the saddle point, the geometric phase satisfies $\Delta S_{\text{geom}}=0$, and the summation over winding sectors produces a geometric series that diverges at the corresponding energies. The contribution of multiple windings suppresses transverse fluctuations, localizing the twistor moduli near the saddle point. The stability of this structure is maintained by a topological constraint: when a branch point approaches the integration contour, it encounters its conjugate image, leading to a pinch singularity that preserves the winding sector.
In this solution of the loop equations, the Master Field is realized as a trajectory in twistor space, analogous to spectral curves in Seiberg–Witten theory. Because the bulk geometry is determined by the boundary data rather than fluctuating dynamically, this construction does not involve a propagating graviton sector and instead provides a geometric realization of gauge holography.
In this framework, the quantum dynamics of planar QCD, encoded in the loop equations and reproduced by planar diagrams, are realized as classical dynamics of the dual geometric variables. The fermion determinant on the Hodge-dual minimal surface reproduces the quantum planar expansion, while the underlying geometry remains classical, reflecting the Master Field nature of the large-$N_{c}$ limit.
In this way, the QCD mass spectrum was computed by solving a problem of classical geometry.
@@ Acknowledgements
The author thanks N. Arkani-Hamed for the invitation to present this work at the IAS Particle Physics (Pizza) Seminar, and the participants of the seminar for detailed and stimulating discussions.
The author is also grateful to E. Witten for discussions that clarified the relation between this twistor-geometric construction and the QCD string observed in lattice studies, as well as for comments on its topological aspects.
@@
During the preparation of this manuscript, the author used the generative AI systems Gemini 3 Flash and ChatGPT 5.3 for technical assistance in editing and presentation. The tools were used for correcting typographical and grammatical errors, improving the clarity of exposition, and standardizing notation and LaTeX formatting.
All scientific content, derivations, and conclusions were developed and verified by the author, who takes full responsibility for the manuscript.
@@ Appendix C Explicit invariant tensor solutions for the MLE
Table 2: Linear algebra diagnostics for the exact Vector Momentum Loop Equation. Here $\Delta\text{rank}:=\text{rank}([A_{n}|b_{n}])-\text{rank}(A_{n})$. A strict positivity, $\Delta\text{rank}>0$, proves the system is mathematically inconsistent.
@@ C.1 Solving for the lowest terms $W^{(n<8)}$
The tensor $\mathcal{W}^{(6)}_{\mu_{1}\dots\mu_{6}}$ is completely determined by the geometric stress cascading from the left-hand side loop derivative acting on the lower-order terms, absorbing the kinematic constraints without contradiction. Its explicit expansion in terms of the 15 independent 6-point Kronecker pairings is uniquely fixed by the lower-order parameters $c_{2,1}\equiv 1$, $c_{4,1}$, and $c_{4,2}$.
At $\mathcal{W}^{(8)}$, this linear absorption fails. The expansion truncates due to the irreducible non-linear $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$ stress, which exceeds the tensor rank of the pure cyclic Shuffle Ideal basis, proving the non-analyticity of the functional $\mathcal{W}[P]$.
@@ C.2 Counting of unknowns/equations and Fredholm inconsistency at $\mathcal{W}^{(8)}$
To demonstrate the breakdown of the Taylor-Magnus expansion at the 8th order ($\mathcal{W}^{(8)}$), we cast the Momentum Loop Equation at each order $n$ into a finite linear system. After imposing the closed-loop shuffle-ideal reduction (loop-closing prescription), cyclic symmetry, and parity constraints (all odd orders vanish), the functional matching reduces to:
@@
The total number of unknown variables solved for at order $n$ is therefore $N_{unk}(n)=N_{\mathcal{W}}(n)+N_{free}(<n)+N_{gauge}(n)$. We analyzed the exact linear algebra of these systems using a specialized symbolic script. The matrix ranks and subsystem counts are summarized in Table 2.
@@
At $n=8$, the matrix $A_{8}$ possesses 297 left-null vectors. Projecting the inhomogeneous right-hand side $b_{8}$ (the geometric stress generated by the discrete non-linear cross-terms of $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$) onto these null directions yields explicit, non-zero rational obstructions. For example, selecting the first canonical left-null basis vector $u_{1}$ yields the rational projection:
@@
This establishes a mathematical contradiction. Two critical robustness statements follow directly from this matrix representation:
@@
The full basis construction, shuffle-ideal reduction, and linear-algebra verification—including the generation of the matrices $A_{n}$ and the explicit null-vector projections—are provided in the executable Mathematica notebook [30] (see Data Availability).
Appendix D The Pinch Singularity and the Topological Barrier